OpenStudy (anonymous):

please help with the process of differentiating 4/ 3x^3

OpenStudy (anonymous):

So you want do calculate\[\frac{d}{dx} \frac{4}{3x^{3}}\]?

OpenStudy (anonymous):

ye think so

OpenStudy (anonymous):

Ok.\[\frac{d}{dx} \frac{4}{3x^{3}}=\frac{4}{3} \frac{d}{dx} x^{-3}\]Now you can just use the power rule:\[\frac{d}{dx}x^{a}=a \times x^{a-1}\]where a=-3.

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

ye that makes sense, thanks

OpenStudy (anonymous):

:)